# Scaling RL to Long Videos

### v3

@@ Abstract
Abstract: We introduce a full-stack framework that scales up reasoning in vision-language models (VLMs) to long videos, leveraging reinforcement learning. We address the unique challenges of long video reasoning by integrating three critical components: (1) a large-scale dataset, LongVideo-Reason, comprising 104K long video QA pairs with high-quality reasoning annotations across diverse domains such as sports, games, and vlogs; (2) a two-stage training pipeline that extends VLMs with chain-of-thought supervised fine-tuning (CoT-SFT) and reinforcement learning (RL); and (3) a training infrastructure for long video RL, named Multi-modal Reinforcement Sequence Parallelism (MR-SP), which incorporates sequence parallelism and a vLLM-based engine tailored for long video, using cached video embeddings for efficient rollout and prefilling. In our experiments, LongVILA-R1-7B achieves strong performance on video benchmarks, reaching 65.1% and 71.1% accuracy on VideoMME without and with subtitles, respectively, and consistently outperforming LongVILA-7B across multiple benchmarks. Moreover, LongVILA-R1-7B supports processing up to 8,192 video frames per video, and configurable FPS settings. Notably, our MR-SP system achieves up to 2.1$\times$ speedup on long video RL training. In addition, we release our training system for public availability that supports RL training on various modalities (video, text, and audio), various models (VILA and Qwen series), and even image and video generation models. On a single A100 node (8 GPUs), it supports RL training on hour-long videos (e.g., 3,600 frames). Code and models are available at https://github.com/NVlabs/Long-RL and https://huggingface.co/Efficient-Large-Model/LongVILA-R1-7B.
@@ 1 Introduction
Understanding long videos requires more than simple recognition—it demands reasoning from temporal, spatial, goal-oriented, and narrative perspectives [1]. As illustrated in Figure 2, answering high-level questions often hinges on a model’s ability to integrate clues distributed across time, infer hidden goals or strategies, track entities spatially, and comprehend the evolving plot. For instance, predicting the winner of a football penalty shootout involves assessing emotional cues and tactical behavior (temporal and goal reasoning), while determining the final location of a hidden ball requires precise spatial tracking. Likewise, evaluating a poker player’s decision demands interpreting implicit strategies beyond surface actions (goal reasoning) and understanding a character’s development or match trajectory reflects the need for plot reasoning. These examples underscore that reasoning is indispensable for long video understanding that goes beyond recognition alone.
Figure 1: Training efficiency comparison with MR-SP (SP degree=4) on Qwen2.5-VL-7B and LongVILA-R1-7B and a single node 8$\times$ A100 GPUs. It achieves 2.1$\times$ speed-up and avoids GPU OOM issue on long frames.
@@
Despite the clear importance of reasoning in long video understanding, enabling such capabilities in long video VLMs poses significant challenges [2, 3, 4, 5, 6]. First, the collection of high quality long video reasoning datasets is inherently difficult. Unlike domains such as math or code reasoning, where structured supervision and benchmarks are readily available [7, 8], long video reasoning requires annotating complex temporal dynamics, goals, spatial relations, and narrative elements—often across minutes or hours of footage [9]. This process is labor-intensive and subjective, making large-scale dataset construction slow and costly. Second, the RL training framework for long videos is challenging. Reinforcement learning, a common strategy for aligning models with complex reasoning objectives, is computationally expensive and sample-inefficient [10]. When applied to long videos, RL becomes even more burdensome due to the extended video frames, requiring more memory and longer rollout runtime. These challenges jointly hinder the development of effective long video VLMs with strong reasoning capabilities.
@@
In this work, we introduce LongVILA-R1, a comprehensive framework exploring the reasoning capabilities for long video understandings. Firstly, we strategically construct a high quality dataset with CoT annotations for long video reasoning, named LongVideo-Reason. Leveraging a powerful VLM (NVILA-8B) [11] and a leading open-source reasoning LLM, we develop a dataset comprising 104K high quality Question-Reasoning-Answer pairs for long videos. We use 36K high quality samples for Long-CoT-SFT to initialize the model’s reasoning and instruction-following abilities, and 68K samples with an additional 102K video data [12, 13, 14, 15, 16] for reinforcement learning. This two-stage training combines high quality reasoning annotations with reinforcement learning, enabling LongVILA-R1 to achieve superior and generalized video reasoning. We also manually curate a balanced set of 1K long video samples to build a new benchmark, LongVideo-Reason-eval, that evaluates performance from four perspectives: Temporal, Goal and Purpose, Spatial, and Plot and Narrative, for a comprehensive assessment.
Figure 3: Data Distribution of LongVideo-Reason and total data in the LongVILA-R1 training framework. LongVideo-Reason comprises a total of 18K videos and 104K QAs with reasoning annotations. Additionally, we also include additional 102K QAs from existing works [12, 13, 14, 15, 16] for RL training.
Secondly, we propose a training framework for VLMs to advance long video reasoning. As illustrated in Figure 5, this framework incorporates two stages, i.e., Stage-1: Long CoT-SFT, and Stage-2: RL for long video reasoning. To address the unique challenges of long video RL, including massive visual embeddings, heavy rollouts, and long-context LLM prefilling, we develop an efficient and scalable solution, referred to as Multi-modal Reinforcement Sequence Parallelism (MR-SP). It incorporates a vLLM engine [17] tailored for LongVILA and a caching scheme for video embeddings. The MR-SP system alleviates the problem of intensive memory and facilitates RL training of long video VLMs. As shown in Figure 1, our MR-SP system achieves up to 2.1$\times$ speedup on 512-frame video RL training on 7B models and enables longer training frames without out-of-memory (OOM).
@@
In our experiments, LongVILA-R1-7B demonstrates strong performance on video benchmarks, achieving 65.1% and 71.1% accuracy on VideoMME without and with subtitles, respectively. It consistently outperforms LongVILA-7B across a range of benchmarks, including ActivityNet-QA, LongVideoBench, PerceptionTest, NExT-QA, VNBench, and VideoMME. Additionally, LongVILA-R1 supports processing up to 8,192 video frames per video, and configurable FPS settings. On our LongVideo-Reason-eval benchmark, LongVILA-R1-7B achieves an average accuracy of 72.0%, surpassing open-source models like Video-R1-7B[2], as well as proprietary models such as Gemini-1.5-Pro[18].
Figure 4: Data generation process for the LongVideo-Reason dataset. This process begins with segmenting videos into 10-second clips and generating captions for each clip using NVILA-8B. Then based on the captions of all clips in a video, we generate question-answer pairs that involve reasoning across the content of the whole video, along with the reasoning annotations using a leading open-source reasoning LLM. Reasoning questions are categorized into Temporal, Goal and Purpose, Spatial, and Plot and Narrative. Finally, the reasoning annotations are reformatted for conciseness and alignment with video details. We present a more detailed figure of data generation process in Figure 9 in the appendix.
@@ 2 Related Work
The field of multi-modal reasoning has advanced significantly, particularly in Vision-Language Models (VLMs). GPT-4o [19] improves visual understanding through enhanced reasoning, while Gemini-1.5-Pro [18] extends context length to 1 million tokens, achieving state-of-the-art performance in VideoMME [20]. Following the substantial progress of architecture and training algorithms in open-source VLMs [21, 11, 22, 23, 24], multi-modal reasoning has been further explored in works including LMM-R1 [25] which employs a two-stage training strategy, Vision-R1 [26] that addresses post-cold-start overthinking, and Video-R1 [2] which enhances RL for video via T-GRPO in 16 frames. However, these approaches primarily focus on single images or short videos, and long video reasoning still poses great challenges.
@@
Training with long contexts often exceeds the memory capacity of a single device [27], necessitating efficient distribution strategies. Sequence parallelism (SP) has become a widely adopted solution  [28, 29, 30, 31, 32]. For example, ring-based systems like LightSeq [33] and Ring Attention [28] use point-to-point (P2P) communication, while DeepSpeed-Ulysses [29] employs all-to-all (A2A) primitives to optimize attention computations. Additionally, USP and LoongTrain  [31, 32] were introduced to integrate Ring-style SP and Ulysses SP. LongVILA [4] further proposed multi-modal SP (MM-SP), enabling vision-language models to handle long-context inputs. However, multi-modal reinforcement learning introduces additional challenges, as it requires extensive sampling from long, mixed-token sequences [10], particularly in complex group optimization tasks [7].
@@
Reinforcement Learning (RL) has become a key strategy for enhancing Large Language Models (LLMs), particularly through Reinforcement Learning with Human Feedback (RLHF) [34] or Direct Preference Optimization [35], which aligns model outputs with human preferences. Recent advancements demonstrate that RL significantly improves LLM reasoning abilities. For example, DeepSeek-R1 [36] utilizes the Group Relative Policy Optimization (GRPO) algorithm [37], integrating group-based sampling and rule-based rewards. On the other hand, RL [38] poses a unique challenge of heavy computational cost, especially in multi-modal settings. To address this problem, HybridFlow [10] is introduced, leveraging Ray [39] for efficient data flow and vLLM [17] for faster sampling. Nevertheless, this remains a bottleneck when processing long video sequences, with group-based sampling constrained by the high computational cost of long-context sampling and visual encoding. In this work, we propose MR-SP that provides up to 2.1$\times$ speedup for long video RL training.
@@ 3.1 Overview of Data Curation
We first curate 18K long videos from the Shot2Story dataset [40] (Figure 3, left). We further incorporated 2k additional 4K-resolution videos spanning scenarios such as autonomous driving, video games, household robotics, and wildlife. We then apply a high quality automated annotation pipeline for CoT as detailed in Section 3.2, and end up with a total of 104K Question-Reasoning-Answer pairs where each sample, based on the type of question it is reasoning about, can be categorized into Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning (Figure 3, middle). This dataset is designed to support various types of long-video reasoning tasks comprehensively.
Given the sensitivity of GRPO to batch sampling [41, 25], a data filtering approach is adopted. Specifically, we employ a test-scaling method where LongVILA [4] performs inference 10 times on the original datasets. Questions consistently answered correctly or incorrectly are labeled as easy or hard while those inducing diverse predictions are labeled as medium. We filter out too easy and too hard questions, and use remaining questions for training. The reason is that GRPO expects different rollouts of each sample to be diverse in order to have meaningful advantages, and the gradient vanishes if all the rollouts predict correct or incorrect answers. The COT-SFT subset (36K) features high quality CoT reasoning processes formatted in a standard <think></think><answer></answer> structure, providing abundant resources for warm-up training during Stage 1 of the model’s reasoning capabilities. Meanwhile, the RL subset contains 68K challenging long-video Q&A, which is leveraged in Stage 2 for scaling reasoning through reinforcement learning. To further enhance RL scaling, we incorporate an additional 102K high quality open-source videos (Figure 3, right) from other datasets [12, 13, 14, 15, 16]. This combination improves the model’s generalization.
Figure 5: The LongVILA-R1 training pipeline. LongVILA-R1 builds upon the base training pipeline for LongVILA. MM-SP is further employed for SFT on long video understanding tasks with long CoT. Then, reinforcement scaling learning is conducted through Multi-modal Reinforcement Sequential Parallelism (MR-SP).
@@ 3.2 Long-Video Reasoning Generation
We introduce an automated annotation pipeline (Figure 4) that generates high quality Question-Reasoning-Answer pairs from long videos. This process begins by segmenting videos into short clips (~10 seconds each), each of which is annotated using the NVILA-8B [11] model to provide descriptive captions. For the spatial reasoning category, We employed VILA-HD [24] to generate object bounding boxes in video frames and constructed spatial reasoning QAs based on these boxes and corresponding captions. Leveraging the breakthrough in text-based reasoning, we then deploy a leading open-source reasoning LLM, DeepSeek-R1-671B [36], provide the captions of all the clips in each video, and prompt it to generate diverse types of Question-Reasoning-Answer pairs that involve reasoning over the content across the whole video.
@@ 4 LongVILA Training Pipeline
As shown in Figure 5, there are two extended training stages in LongVILA-R1, i.e., (1) warm-up for long video reasoning, utilizing 36K data with high quality CoT for SFT on the MM-SP system [4]; (2) reinforcement learning with dense frames from long videos.
@@ 4.1 Long Video CoT Supervised Fine-Tuning
Utilizing 104K high quality question-reasoning-answer pairs, we apply the data filtering method described in Section 3.1 to select 36K examples for long CoT-SFT, serving as a warm-up phase for subsequent RL. This stage equips the model with fundamental reasoning abilities and instruction-following skills for long video scenarios. To efficiently perform SFT on hundreds of frames, we adopt the MM-SP [4] training system from LongVILA. As demonstrated in Section 6.2, SFT solely on our LongVideo-Reason dataset also effectively improves the model with foundational reasoning capabilities.
@@ 4.2 GRPO for Long Video
Building on the advancements of the GRPO [7] algorithm and prior explorations of multi-modal reasoning training [2, 25], we adhere to the standard GRPO framework to train our model. For each given question $q$, the policy model generates a group of candidate responses $\{o_{1},o_{2},...,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$, accompanied by their corresponding rewards $\{r_{1},r_{2},...,r_{G}\}$, which are computed based on rule-based reward functions (format/accuracy). The model $\pi_{\theta}$ is subsequently optimized by maximizing the following objective function:
@@
Considering the sensitivity of GRPO to sampling during training [41], we use the 68K filtered data for reinforcement learning as described in Section 3.1. Additionally, an extra 102K samples from other datasets [12, 13, 14, 15, 16] are incorporated to scale up the RL. This approach aims to guide the model in freely exploring and developing more effective and generalized reasoning strategies.
@@ 5 Multi-modal Reinforcement SP
Existing RL frameworks for VLMs, such as R1-V [38] and EasyR1 [42], are not designed for long videos which present unique challenges due to their high token volume. To address this, we introduce Multi-modal Reinforcement Sequence Parallelism (MR-SP), a framework for efficient RL training on long videos. MR-SP leverages sequence parallelism in both rollout and pre-filling stages, enabling long videos in RL, with reduced overhead.

### v4

@@ Abstract
[…]
00footnotetext: Equal contribution    Core contribution
@@ 1 Introduction
Understanding long videos requires more than simple recognition—it demands reasoning from temporal, spatial, goal-oriented, and narrative perspectives zhang2024long . As illustrated in Figure 1, answering high-level questions often hinges on a model’s ability to integrate clues distributed across time, infer hidden goals or strategies, track entities spatially, and comprehend the evolving plot. For instance, predicting the winner of a football penalty shootout involves assessing emotional cues and tactical behavior (temporal and goal reasoning), while determining the final location of a hidden ball requires precise spatial tracking. Likewise, evaluating a poker player’s decision demands interpreting implicit strategies beyond surface actions (goal reasoning) and understanding a character’s development or match trajectory reflects the need for plot reasoning. These examples underscore that reasoning is indispensable for long video understanding that goes beyond recognition alone. Despite the clear importance of reasoning in long video understanding, enabling such capabilities in long video VLMs poses significant challenges feng2025video ; weng2024longvlm ; chen2024longvila ; guo2025fila ; shen2025long . First, the collection of high quality long video reasoning datasets is inherently difficult. Unlike domains such as math or code reasoning, where structured supervision and benchmarks are readily available shao2024deepseekmath ; liu2024codemind , long video reasoning requires annotating complex temporal dynamics, goals, spatial relations, and narrative elements—often across minutes or hours of footage han2024videoespresso . This process is labor-intensive and subjective, making large-scale dataset construction slow and costly. Second, the RL training framework for long videos is challenging. Reinforcement learning, a common strategy for aligning models with complex reasoning objectives, is computationally expensive and sample-inefficient sheng2024hybridflow . When applied to long videos, RL becomes even more burdensome due to the extended video frames, requiring more memory and longer rollout runtime. These challenges jointly hinder the development of effective long video VLMs with strong reasoning capabilities.
@@
In this work, we introduce LongVILA-R1, a comprehensive framework exploring the reasoning capabilities for long video understandings. Firstly, we strategically construct a high quality dataset with CoT annotations for long video reasoning, named LongVideo-Reason. Leveraging a powerful VLM (NVILA-8B) liu2024nvila and a leading open-source reasoning LLM, we develop a dataset comprising 104K high quality Question-Reasoning-Answer pairs for long videos. We use 36K high quality samples for Long-CoT-SFT to initialize the model’s reasoning and instruction-following abilities, and 68K samples with an additional 102K video data llava-video ; next-qa ; perpceptiontest ; clevr ; star for reinforcement learning. This two-stage training combines high quality reasoning annotations with reinforcement learning, enabling LongVILA-R1 to achieve superior and generalized video reasoning. We also manually curate a balanced set of 1K long video samples to build a new benchmark, LongVideo-Reason-eval, that evaluates performance from four perspectives: Temporal, Goal and Purpose, Spatial, and Plot and Narrative, for a comprehensive assessment.
[…]
@@
In our experiments, LongVILA-R1-7B demonstrates strong performance on video benchmarks, achieving 65.1% and 71.1% accuracy on VideoMME without and with subtitles, respectively. It consistently outperforms LongVILA-7B across a range of benchmarks, including ActivityNet-QA, LongVideoBench, PerceptionTest, NExT-QA, VNBench, and VideoMME. Additionally, LongVILA-R1 supports processing up to 8,192 video frames per video, and configurable FPS settings. On our LongVideo-Reason-eval benchmark, LongVILA‑R1‑7B achieves an average accuracy of 72.0%, surpassing Video‑R1‑7Bfeng2025video , as well as proprietary models such as Gemini‑1.5‑Proteam2024gemini .
@@ 2 Related Work
The field of multi-modal reasoning has advanced significantly, particularly in Vision-Language Models (VLMs). GPT-4o gpt4o improves visual understanding through enhanced reasoning, while Gemini-1.5-Pro team2024gemini extends context length to 1 million tokens, achieving state-of-the-art performance in VideoMME video-mme . Following the substantial progress of architecture and training algorithms in open-source VLMs zhu2025internvl3 ; liu2024nvila ; yang2024qwen2 ; lin2024vila ; vila-hd , multi-modal reasoning has been further explored in works including LMM-R1 peng2025lmm which employs a two-stage training strategy, Vision-R1 huang2025vision that addresses post-cold-start overthinking, and Video-R1 feng2025video which enhances RL for video via T-GRPO in 16 frames. However, these approaches primarily focus on single images or short videos, and long video reasoning still poses great challenges.
Figure 3: Data Distribution of LongVideo-Reason and total data in the LongVILA-R1 training framework. LongVideo-Reason comprises a total of 18K videos and 104K QAs with reasoning annotations. Additionally, we include 102K QAs from existing works llava-video ; next-qa ; perpceptiontest ; clevr ; star .
@@
Training with long contexts often exceeds the memory capacity of a single device longlora , necessitating efficient distribution strategies. Sequence parallelism (SP) has become a widely adopted solution  liu2023ring ; jacobs2023deepspeed ; InternEvo ; USP ; LoongTrain . For example, ring-based systems like LightSeq li2023lightseq and Ring Attention liu2023ring use point-to-point (P2P) communication, while DeepSpeed-Ulysses jacobs2023deepspeed employs all-to-all (A2A) primitives to optimize attention computations. Additionally, USP and LoongTrain  USP ; LoongTrain were introduced to integrate Ring-style SP and Ulysses SP. LongVILA chen2024longvila further proposed multi-modal SP (MM-SP), enabling vision-language models to handle long-context inputs. However, multi-modal reinforcement learning introduces additional challenges, as it requires extensive sampling from long, mixed-token sequences sheng2024hybridflow , particularly in complex group optimization tasks shao2024deepseekmath .
@@
[…]
@@ 3.1 Overview of Data Curation
We first curate 18K long videos from the Shot2Story dataset han2023shot2story20k (Figure 3, left). We further incorporated 2k additional 4K-resolution videos spanning scenarios such as autonomous driving, video games, household robotics, and wildlife. We then apply a high quality automated annotation pipeline for CoT as detailed in Section 3.2, and end up with a total of 104K Question-Reasoning-Answer pairs where each sample, based on the type of question it is reasoning about, can be categorized into Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning (Figure 3, middle). This dataset is designed to support various types of long-video reasoning tasks comprehensively.
Given the sensitivity of GRPO to batch sampling team2025kimi ; peng2025lmm , a data filtering approach is adopted. Specifically, we employ a test-scaling method where LongVILA chen2024longvila performs inference 10 times on the original datasets. Questions consistently answered correctly or incorrectly are labeled as easy or hard while those inducing diverse predictions are labeled as medium. We filter out too easy and too hard questions, and use remaining questions for training. The reason is that GRPO expects different rollouts of each sample to be diverse in order to have meaningful advantages, and the gradient vanishes if all the rollouts predict correct or incorrect answers. The COT-SFT subset (36K) features high quality CoT reasoning processes formatted in a standard <think></think><answer></answer> structure, providing abundant resources for warm-up training during Stage 1 of the model’s reasoning capabilities. Meanwhile, the RL subset contains 68K challenging long-video Q&A, which is leveraged in Stage 2 for scaling reasoning through reinforcement learning. To further enhance RL scaling, we incorporate an additional 102K high quality open-source videos (Figure 3, right) from other datasets llava-video ; next-qa ; perpceptiontest ; clevr ; star . This combination improves the model’s generalization.
Figure 4: Data generation process for the LongVideo-Reason dataset. This process begins with segmenting videos into 10-second clips and generating captions for each clip using NVILA-8B. Then based on the captions of all clips in a video, we generate question-answer pairs that involve reasoning across the content of the whole video, along with the reasoning annotations using a leading open-source reasoning LLM. Reasoning questions are categorized into Temporal, Goal and Purpose, Spatial, and Plot and Narrative. Finally, the reasoning annotations are reformatted for conciseness and alignment with video details. We present a more detailed figure of data generation process in Figure 9.
@@ 3.2 Long-Video Reasoning Generation
We introduce an automated annotation pipeline (Figure 4) that generates high quality Question-Reasoning-Answer pairs from long videos. This process begins by segmenting videos into short clips (~10 seconds each), each of which is annotated using the NVILA-8B liu2024nvila model to provide descriptive captions. For the spatial reasoning category, We employed VILA-HD vila-hd to generate object bounding boxes in video frames and constructed spatial reasoning QAs based on these boxes and corresponding captions. Leveraging the breakthrough in text-based reasoning, we then deploy a leading open-source reasoning LLM, DeepSeek-R1-671B guo2025deepseek , provide the captions of all the clips in each video, and prompt it to generate diverse types of Question-Reasoning-Answer pairs that involve reasoning over the content across the whole video.
@@ 4 LongVILA Training Pipeline
As shown in Figure 5, there are two extended training stages in LongVILA-R1, i.e., (1) warm-up for long video reasoning, utilizing 36K data with high quality CoT for SFT on the MM-SP system chen2024longvila ; (2) reinforcement learning with dense frames from long videos.
@@ 4.1 Long Video CoT Supervised Fine-Tuning
Utilizing 104K high quality question-reasoning-answer pairs, we apply the data filtering method described in Section 3.1 to select 36K examples for long CoT-SFT, serving as a warm-up phase for subsequent RL. This stage equips the model with fundamental reasoning abilities and instruction-following skills for long video scenarios. To efficiently perform SFT on hundreds of frames, we adopt the MM-SP chen2024longvila training system from LongVILA. As demonstrated in Section 6.2, SFT solely on our LongVideo-Reason dataset also effectively improves the model with reasoning capabilities.
@@ 4.2 GRPO for Long Video
Building on the advancements of the GRPO shao2024deepseekmath algorithm and prior explorations of multi-modal reasoning training feng2025video ; peng2025lmm , we adhere to the standard GRPO framework to train our model. For each given question $q$, the policy model generates a group of candidate responses $\{o_{1},o_{2},...,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$, accompanied by their corresponding rewards $\{r_{1},r_{2},...,r_{G}\}$, which are computed based on rule-based reward functions (format/accuracy). The model $\pi_{\theta}$ is subsequently optimized by maximizing the following objective function:
@@
Figure 5: The LongVILA-R1 training pipeline. LongVILA-R1 builds upon the base training pipeline for LongVILA. $M$-SP is further employed for SFT on long video understanding tasks with long CoT. Then, reinforcement scaling learning is conducted through Multi-modal Reinforcement Sequential Parallelism (MR-SP).
@@
Considering the sensitivity of GRPO to sampling during training team2025kimi , we use the 68K filtered data for reinforcement learning as described in Section 3.1. Additionally, an extra 102K samples from other datasets llava-video ; next-qa ; perpceptiontest ; clevr ; star are incorporated to scale up the RL. This approach aims to guide the model in freely exploring and developing more effective and generalized reasoning strategies.
@@ 5 Multi-modal Reinforcement SP
Existing RL frameworks for VLMs, such as R1-V chen2025r1v and EasyR1 EasyR1 , are not designed for long videos which present unique challenges due to their high token volume. To address this, we introduce Multi-modal Reinforcement Sequence Parallelism (MR-SP), a framework for efficient RL training on long videos. MR-SP leverages sequence parallelism in both rollout and pre-filling stages, enabling long videos in RL, with reduced overhead. We show the training curve with MR-SP in Figure 8.